OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (jdoe0001):

if you post specifics, rather then general material, we could tell if we could or not

OpenStudy (abbycross167):

Please help me to simplify this rational expression - x+1 3x+3 ----- / ----- 4 16 **Note the / means to divide

OpenStudy (phi):

what is the first step when dividing by a fraction ? change it by "flipping" and multiplying, right ?

OpenStudy (jdoe0001):

\(\cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\qquad thus \\ \quad \\ \cfrac{\frac{x+1}{4}}{\frac{3x+3}{16}}\implies \cfrac{x+1}{4}\cdot \cfrac{16}{3x+3}\implies \cfrac{\cancel{x+1}}{\cancel{4}}\cdot \cfrac{\cancel{16}}{3\cancel{(x+1)}}\implies ?\)